### v4

# Proposal for simulating quantum spin models with Dzyaloshinskii-Moriya interaction using Rydberg atoms, and construction of asymptotic quantum many-body scar states

Abstract
We have developed a method to simulate quantum spin models with Dzyaloshinskii-Moriya interaction (DMI) using Rydberg atom quantum simulators. Our approach involves a two-photon Raman transition and a transformation to the spin-rotating frame, both of which are feasible with current experimental techniques. As a model that can be simulated in our setup but not in solid-state systems, we consider an $S=1/2$ spin chain with a Hamiltonian consisting of the DMI and Zeeman energy. We study the magnetization curve in the ground state of this model and quench dynamics. Further, we show the existence of quantum many-body scar states and asymptotic quantum many-body scar states. The observed nonergodicity in this model demonstrates the importance of the highly tunable DMI that can be realized by the proposed quantum simulator.
I Introduction
In recent years, quantum simulators have become powerful tools for exploring various quantum many-body phenomena. These simulators, which include ultracold gases Bloch2008 ; Gross2017 ; Schafer2020 , Rydberg atoms Browaeys2020 ; Morgado2021 , trapped ions Blatt2012 ; Monroe2021 , and superconducting qubits Wendin2017 ; Kjaergaard2020 ; Altman2022 , offer high controllability, allowing researchers to realize different types of Hamiltonians in experiments. Rydberg atom quantum simulators, in particular, have attracted much attention due to their ability to create quantum spin models. By leveraging the strong dipole-dipole interaction between highly excited atoms and optical tweezers in a programmable array, these simulators have successfully realized a range of models, such as Ising Labuhn2016 ; Zeiher2017 ; Bernien2017 ; Leseleuc2018 ; Lienhard2018 ; Guardado-Sanchez2018 ; Keesling2019 ; Ebadi2021 ; Semeghini2021 ; Bluvstein2021 ; Scholl2021 ; Bharti2022a , XY Orioli2018 ; Leseleuc2019 ; Chen2023 , XXZ Signoles2021 ; Geier2021 ; Scholl2022 ; Franz2022a , and XYZ Geier2021 ; Steinert2023 models. This versatility makes them valuable for investigating issues in statistical mechanics, condensed matter physics, and quantum computing. Notable achievements in Rydberg atom quantum simulators include the discovery of quantum many-body scar states Bernien2017 ; Turner2018 ; Turner2018_2 , observations of topological edge states Leseleuc2019 , and realizations of quantum spin liquid states Semeghini2021 . Furthermore, researchers have successfully implemented several gate operations Levine2018 ; Levine2019 ; Madjarov2020 ; Bluvstein2022 ; Graham2022 ; Jenkins2022 ; Ma2022 ; Wu2022 ; Chew2022 , demonstrating the potential of these simulators for advancing quantum technologies. These developments highlight the promising role of quantum simulators in advancing our understanding of complex quantum systems and their applications in various scientific and technological domains.
In solid-state physics, the Dzyaloshinskii-Moriya interaction (DMI) Dzyaloshinsky1958 ; Moriya1960 is crucial in understanding magnetism in crystals without space-inversion symmetry. This interaction leads to nontrivial magnetic structures in materials with structural chirality, known as chiral magnets. The classical effects of DMI include the formation of chiral soliton lattices Togawa2012 ; Kishine2015 ; Togawa2016 and skyrmions Muhlbauer2009 ; Yu2010 ; Nagaosa2013 . Moreover, the role of DMI in quantum phenomena has been extensively studied Oshikawa1997 ; Affleck1999 ; Takashima2016 ; Kodama_thesis ; Kodama2023 . One of the most conspicuous quantum effects in magnets is the spin parity effect. This effect refers to the phenomenon in which properties differ significantly depending on whether the spin is half-integer or integer. Recent theoretical work by Kodama et al. has shed light on the spin parity effects in one-dimensional chiral magnets, particularly when the DMI exceeds the exchange interaction Kodama_thesis ; Kodama2023 . However, experimental observation of these effects is challenging due to the relatively small magnitude of the DMI compared to the exchange interaction in typical crystal systems. A promising candidate for quantum simulations of the DMI is neutral cold atoms in optical lattices with synthetic spin-orbit coupling (SOC) Dalibard2011 ; Goldman2014 ; Zhai2015 . Previous theoretical studies have shown that an effective Hamiltonian in the strongly correlated regimes becomes the spin Hamiltonian with the DMI Cole2012 ; Radic2012 ; Cai2012 ; Peotta2014 ; Gong2015 ; Goldman2023a . However, this synthetic SOC induces heating, which makes the low-energy physics and long-time evolution inaccessible Wang2012 ; Huang2016 ; Wu2016 ; Sun2018 .
The purpose of this paper is two-fold. First, we propose a method to simulate quantum spin models with DMI in Rydberg atom quantum simulators, in a way accessible with current experimental techniques. Second, we explore the nontrivial properties of a quantum spin model that exemplifies the importance of our quantum simulator.
We can obtain the effective DMI in the spin-rotating frame using two-photon Raman transition and unitary transformation. In our scheme, we can tune the ratio between the DMI and the exchange interaction from zero to infinity. This tunability allows us to realize the Hamiltonian with only DMI and magnetic field term, called the DH model Kodama_thesis ; Kodama2023 . We investigate the ground-state properties and nonequilibrium dynamics of the DH model using matrix-product state methods Schollwock2011 ; Paeckel2019 . We thereby show that the time-averaged magnetization continuously changes as a function of the magnetic field in the fully quantum case, whereas it is discontinuous in the classical case. We show that the DH model has QMBS states under periodic and open boundary conditions and argue that they are responsible for the slow thermalization after a quench from a particular initial state. We also show that the DH model has asymptotic quantum many-body scar (AQMBS) states, which are not eigenstates of the Hamiltonian but whose energy variance vanishes in the thermodynamic limit Gotta2023 . We analytically construct a series of AQMBS states and find the slow thermalization after a quench from such states.
This paper is organized as follows: In Sec. II, we explain our proposal for simulating spin models with DMI. In Secs. III and IV, we analyze the ground-state and sweep dynamics of the DH model, respectively. In Secs. V and VI, we show the existence of the QMBS states and AQMBS states in the DH model, respectively. In Sec. VII, we summarize our results.
II Experimental proposal for simulating Dzyaloshinskii-Moriya interaction
[…]
$\hat{H}_{\rm XXZ}$ $=J\sum_{j=1}^{M-1}(\hat{S}_{j}^{x}\hat{S}_{j+1}^{x}+\hat{S}_{j}^{y}\hat{S}_{j+1}^{y}+\delta\hat{S}_{j}^{z}\hat{S}_{j+1}^{z}),$ (1)
where $M$ is the number of spins, $\hat{S}_{j}^{\mu}$$(\mu=x,y,z)$ is the spin-1/2 operator at site $j$, $J$ is the exchange interaction energy, which originates from the van der Waals interaction between the Rydberg atoms, and $\delta$ is the anisotropy parameter. We consider only the nearest-neighbor (NN) part of the interaction because the magnitude of the next NN interaction is $1/64$ that of the NN interaction note_interaction . See Appendix A for a detailed derivation of the XXZ Hamiltonian from the dipole-dipole interaction between the Rydberg atoms.
[…]
To generate the DMI in this system, we irradiate two linearly polarized Raman lasers to the Rydberg atoms [See Fig. 1 (a)]. One laser propagates along the $x$ axis with Rabi frequency $\Omega_{1}$ and wave number $\bm{k}_{1}=k_{1}\bm{e}_{x}$, and the other propagates along the $y$ axis with Rabi frequency $\Omega_{2}$ and wave number $\bm{k}_{2}=k_{2}\bm{e}_{y}$. We denote the position of the $j$th atom as $\bm{R}_{j}\equiv dj(\cos\theta,0,\sin\theta)$ , where $\theta$ is the angle between the chain and the $x$ axis shown in Fig. 1 (a). We can express the effective Hamiltonian for the $j$th atom as
$\hat{h}_{j}$ $=-\hbar\Omega_{\rm eff}[\cos(qj)\hat{S}_{j}^{x}+\sin(qj)\hat{S}_{j}^{y}]-\hbar\tilde{\Delta}\hat{S}_{j}^{z}.$ (2)
Here, $\Omega_{\rm eff}\equiv\Omega_{1}\Omega_{2}/(2\Delta)$ is the effective Rabi frequency with $\Delta\equiv(\Delta_{1}+\Delta_{2})/2$. The one-photon detuning $\hbar\Delta_{i}\equiv\hbar\omega_{i}-(E_{n_{i}S}-E_{P})\;(i=1,2)$ is written in terms of the frequency of the laser $i$ ($\omega_{i}$) and the energies of $\left|n_{i}S_{1/2}\right\rangle$ and $\left|6P_{3/2}\right\rangle$ states ($E_{n_{i}S_{1/2}}$ and $E_{P}$). We also define $q\equiv k_{1}d\cos\theta$ and the two-photon detuning including the ac-Stark shift $\tilde{\Delta}\equiv-[\Delta_{1}-\Delta_{2}+(\Omega_{1}^{2}-\Omega_{2}^{2})/(4\Delta)]$. See Appendix B for a detailed derivation of the rotating transverse field term. The total Hamiltonian in the spin-laboratory frame note1 is given by $\hat{H}_{\text{s-lab}}\equiv\hat{H}_{\rm XXZ}+\sum_{j}\hat{h}_{j}$.
We then move to the spin-rotating frame. The unitary transformation with the operator Perk1976 ; Calvo1981 ; Oshikawa1997 ; Affleck1999 ; Shekhtman1992 ; Nikuni1993
$\hat{U}_{\text{s-rot}}\equiv\prod_{j=1}^{M}e^{-iqj\hat{S}_{j}^{z}},$ (3)
yields the Hamiltonian in the spin-rotating frame:
$\hat{H}_{\text{s-rot}}$ $\equiv\hat{U}^{\dagger}_{\text{s-rot}}\hat{H}_{\text{s-lab}}\hat{U}_{\text{s-rot}}$
$=J\cos q\sum_{j=1}^{M-1}(\hat{S}_{j}^{x}\hat{S}_{j+1}^{x}+\hat{S}_{j}^{y}\hat{S}_{j+1}^{y})$
$-J\sin q\sum_{j=1}^{M-1}(\hat{S}_{j}^{x}\hat{S}_{j+1}^{y}-\hat{S}_{j}^{y}\hat{S}_{j+1}^{x})$
$+J\delta\sum_{j=1}^{M-1}\hat{S}_{j}^{z}\hat{S}_{j+1}^{z}-\hbar\Omega_{\rm eff}\sum_{j=1}^{M}\hat{S}_{j}^{x}-\hbar\tilde{\Delta}\sum_{j=1}^{M}\hat{S}_{j}^{z}.$ (4)
In the new frame, the DMI appears in the third line of Eq. (4). This expression implies that we can simulate the Hamiltonian with the XXZ and DMI terms under the uniform magnetic field by the XXZ Hamiltonian with the rotating transverse field. In Appendices C and D, we discuss the relation between the spin-laboratory and spin-rotating frames and extension to two-dimensional systems, respectively.
In the following, we discuss the tunability of the Hamiltonian. The ratio between the XY interaction and DMI is given by $\tan(k_{1}d\cos\theta)$. We can tune $\theta$ by arranging the Rydberg atoms in appropriate positions using the optical tweezers [see Fig. 1 (a)]. Thereby the strong DMI regime [$|\tan(k_{1}d\cos\theta)|>1$] can be realized, which is practically inaccessible in solid-state systems. The XY interaction term vanishes under the condition $k_{1}d\cos\theta=\pi/2$, which can be realized when $k_{1}=2\pi/(1015~{}{\rm nm})$, $d=5~{}\mu{\rm m}$, and $\theta\simeq 87.1^{\circ}$. We can also vary the anisotropy parameter $\delta$ and the detunings by choosing different combinations of Rydberg states and applying additional static electric and magnetic fields Whitlock2017 ; Leseleuc2018 ; Weber2017 ; Sibalic2017 ; Robertson2021 . This tunability offers the potential to realize the DH model Kodama_thesis ; Kodama2023 in the Rydberg atom quantum simulators:
$\hat{H}_{\rm DH}=D\sum_{j=1}^{M-1}(\hat{S}_{j}^{x}\hat{S}_{j+1}^{y}-\hat{S}_{j}^{y}\hat{S}_{j+1}^{x})-h^{x}\sum_{j=1}^{M}\hat{S}_{j}^{x},$ (5)
where $D\equiv-J$ is the magnitude of the DMI and $h^{x}\equiv\hbar\Omega_{\rm eff}$ is the magnetic field.
We remark that the DMI can be expressed as the Peierls phase because the first two terms in Eq. (4) reduces to $e^{-iq}\hat{S}_{j}^{+}\hat{S}_{j+1}^{-}+e^{+iq}\hat{S}_{j}^{-}\hat{S}_{j+1}^{+}$. Recently, the methods for controlling the Peierls phase in the Rydberg atom quantum simulators have been studied experimentally Lienhard2020 and theoretically Yang2022 ; Wu2022_2 ; Perciavalle2023 ; Poon2024 .
III Ground state properties of the DH model
In this section, we show the ground state properties of the DH model based on the density matrix renormalization group (DMRG) calculations White1992 ; White1993 . Some of the results have been reported in Ref. Kodama_thesis , but we reproduce them here for completeness.
Figure 2 shows the magnetization curve of the DH model (5) for various system sizes. We see some steep changes in the magnetization curve. Magnetization curves in classical chiral magnets exhibit similar behavior, which we attribute to level crossings between states with different winding numbers Kishine2014 . Although the winding number is not well defined in quantum systems, the soliton number operator
$\hat{N}_{\rm sol}\equiv\sum_{j=1}^{M-1}\left(\frac{1}{4}-\hat{S}_{j}^{x}\hat{S}_{j+1}^{x}\right),$ (6)
plays a similar role Kishine2014 ; Kodama_thesis ; Kodama2023 . We plot the ground state expectation value of $\hat{N}_{\rm sol}$ as a function of the magnetic field in Fig. 3. The behavior is consistent with the magnetization curve shown in Fig. 2. The changes in the expectation value of the soliton number $\langle\hat{N}_{\rm sol}\rangle$ by $1$ correspond to the sharp changes in the magnetization. The local spin density exhibits similar behavior as shown in Fig. 4, where we find a helical structure of the spin in low-field regions. In high-field regions, the spin density is almost uniform except near the edges of the system. When we lower the magnetic field, the solitons enter the system, and the expectation value of $\hat{S}_{\rm tot}^{x}$ reduces.
Figure 2: Magnetization curve of the DH model (5) in the ground states for various system sizes. (a) $M=16$ (b) $M=20$ (c) $M=30$ (d) $M=40$ (e) $M=50$ (f) $M=100$.
Figure 3: Ground state soliton number of the DH model (5) as a function of $h^{x}$ for various system sizes. (a) $M=16$ (b) $M=20$ (c) $M=30$ (d) $M=40$ (e) $M=50$ (f) $M=100$.
Figure 4: Expectation value of the local spin $S_{j}^{x}$ (red circle) and $S_{j}^{y}$ (blue triangle) of the DH model (5) in the ground states for $M=100$. (a) $h^{x}=0.1D$ (b) $h^{x}=0.4D$ (c) $h^{x}=0.6D$ (d) $h^{x}=0.78D$ (e) $h^{x}=0.81D$ (f) $h^{x}=0.85D$. We also show the expectation value of the soliton number operator for each parameter.
IV Sweep dynamics of the DH model
In this section, we calculate the sweep dynamics of the DH model to investigate the quantum nature of the chiral magnets. We consider the DH Hamiltonian with a time-dependent field:
$\hat{H}(t)$ $=D\sum_{j=1}^{M-1}(\hat{S}_{j}^{x}\hat{S}_{j+1}^{y}-\hat{S}_{j}^{y}\hat{S}_{j+1}^{x})-h^{x}(t)\sum_{j=1}^{M}\hat{S}_{j}^{x}.$ (7)
We will discuss the time dependence of $h^{x}(t)$ in the next paragraph. We use the time-evolving block decimation (TEBD) method Vidal2003 ; Vidal2004 with the second-order Trotter decomposition for real-time evolution. We treat the explicit time dependence of the Hamiltonian by the technique proposed in Ref. Hatano2005 . We set the discrete time step to $\Delta t=0.01\hbar/D$, and take the bond dimension to be sufficiently large to avoid artifact of truncation. For comparison, we also calculate the dynamics of the classical DH model by replacing the spin operators with c-numbers and using the standard fourth-order Runge-Kutta method for the time evolution.
We set the initial condition as $\left|\psi(0)\right\rangle=\left|+_{1}+_{2}\ldots+_{M}\right\rangle$, where |±$\left|\pm_{j}\right\rangle\equiv(\left|\uparrow_{j}\right\rangle\pm\left|\downarrow_{j}\right\rangle)/\sqrt{2}$ are the eigenstates of $\hat{S}_{j}^{x}$ with eigenvalues ±$\pm 1/2$. This state is almost the ground state of the DH model for large magnetic-field regimes, as shown in the previous section. We set the magnetic field to $h_{\rm ini}^{x}=1D$ at $t=0$ and start to ramp down the magnetic field $h^{x}$ linearly to the final value $h_{\rm fin}^{x}$ during $0\leq t\leq\tau$, then fix it until $t=\tau+100\hbar/D$ [see the time sequence in Fig. 5 (a)]. Here, we fix the ramp rate to $0.1D^{2}/\hbar$.
Figure 5 (b) shows the time-averaged magnetization in the $x$ direction $\hat{S}_{\rm tot}^{x}\equiv\sum_{j=1}^{M}\hat{S}_{j}^{x}$. We also plot the ground-state magnetization curve obtained by the DMRG calculations. The time-averaged magnetizations do not follow the ground state for both classical and quantum cases. In particular, the time-averaged magnetization of the classical case suddenly changes around $h^{x}_{\rm fin}\sim 0.25D$ [see Figs. 5 (b) and (d)]. This behavior indicates metastability of the uniformly polarized state, which has been discussed in the context of classical chiral magnets Togawa2015 ; Mito2018 ; Shinozaki2018 . The sudden jump is due to the vanishing of the surface energy barrier, which prevents the chiral solitons from entering the bulk from the boundaries of the system. In contrast to the classical case, the time-averaged magnetization changes smoothly as a function of the magnetic field in the quantum case, as shown in Figs. 5 (b) and (c). We attribute this behavior to macroscopic quantum tunneling between states with different winding numbers due to strong quantum fluctuations in $S=1/2$ systems.
We now discuss the experimental observables. In optical tweezer experiments, the $z$-component of the local spin can be extracted from measured recapture probability after depopulating one of the spin components through a short-lived intermediate state Scholl2022 . We can also obtain the $x$- and $y$-components using the quantum-state tomography technique by rotating the measurement basis with a microwave Signoles2021 ; Scholl2022 ; Geier2021 . Since all spin components are observables, the above quantum dynamics can be accessible experimentally.
Figure 5: (a) Time sequence of the magnetic field $h^{x}(t)$. We calculate the time averages in the range $\tau+50\hbar/D\leq t\leq\tau+100\hbar/D$. (b) The $x$ component of the time-averaged magnetization per site for TEBD (blue square) and classical (green triangle) results. The error bars represent the standard deviation of the time average. For reference, we plot the ground-state magnetization calculated by the DMRG method (red circle). (c)-(d) Magnetization of $x$ component vs time for TEBD (c) and classical (d) results. The color bars represent the results for $h^{x}_{\rm fin}=0.9D$ (red) to $h^{x}_{\rm fin}=0.1D$ (blue).
V Quantum many-body scar states of the DH model
In this section, we show that the DH model has QMBS states. The QMBS states Serbyn2021 ; Papic2022 ; Moudgalya2022 ; Chandran2023 have been discussed in the context of thermalization problems in isolated quantum systems DAlessio2016 ; Mori2018 and extensively studied theoretically Choi2019 ; Ho2019 ; Lin2019 ; Schecter2019 ; Iadecola2020 ; Chattopadhyay2020 ; Mark2020 ; Mark2020_2 ; ODea2020 ; Shibata2020 ; Michailidis2020 ; Kuno2020 ; Sugiura2021 ; Tang2022 ; Tamura2022 ; Sanada2023 ; Omiya2023 ; Inversen2023 ; Moudgalya2022a ; Moudgalya2023a and experimentally Bernien2017 ; Bluvstein2021 ; Su2023 . QMBS states refer to special eigenstates in nonintegrable systems that do not satisfy the strong version of the eigenstate thermalization hypothesis (ETH) Deutsch1991 ; Srednicki1994 ; Rigol2008 . If we set the initial condition with a large overlap with a QMBS state, the relaxation to the thermal equilibrium state is slow or does not occur persistently. Under appropriate conditions, the system exhibits persistent oscillations.
In the following, we show that the DH model has the QMBS states based on a restricted spectrum generating algebra (RSGA) Moudgalya2020 . For completeness, we recall the definition of the RSGA of order $1$ before showing the results of the DH model.
Let us consider the Hamiltonian $\hat{H}_{0}$ and the operator $\hat{Q}^{\dagger}$ such that $\hat{Q}^{\dagger}\left|\psi_{0}\right\rangle\not=0$ for some state $\left|\psi_{0}\right\rangle$. We assume that $\hat{H}_{0}$, $\hat{Q}^{\dagger}$, and $\left|\psi_{0}\right\rangle$ satisfy the following conditions:
$\hat{H}_{0}\left|\psi_{0}\right\rangle$ $=E_{0}\left|\psi_{0}\right\rangle,$ (8)
$[\hat{H}_{0},\hat{Q}^{\dagger}]\left|\psi_{0}\right\rangle$ $=\mathcal{E}\hat{Q}^{\dagger}\left|\psi_{0}\right\rangle,$ (9)
$[\,[\hat{H}_{0},\hat{Q}^{\dagger}],\hat{Q}^{\dagger}]$ $=0,$ (10)
where $E_{0}$ and $\mathcal{E}$ are real numbers. These relations are called the RSGA of order 1. If the above relations are satisfied, we can show the following relations:
$\hat{H}_{0}\left|\psi_{n}\right\rangle$ $=(E_{0}+n\mathcal{E})\left|\psi_{n}\right\rangle,$ (11)
$\left|\psi_{n}\right\rangle$ $\equiv(\hat{Q}^{\dagger})^{n}\left|\psi_{0}\right\rangle\quad\text{or }\left|\psi_{n}\right\rangle=0.$ (12)
The state $\left|\psi_{n}\right\rangle$ corresponds to the (unnormalized) QMBS state.
V.1 Periodic boundary case
In this subsection, we show the existence of the QMBS states of the DH model under the periodic boundary conditions. The Hamiltonian is given by
$\hat{H}_{\rm DH}^{\rm PBC}$ $=D\sum_{j=1}^{M}\hat{S}_{j}^{x}(\hat{S}_{j-1}^{z}-\hat{S}_{j+1}^{z})-h^{x}\sum_{j=1}^{M}\hat{S}_{j}^{z},$ (13)
where $\hat{S}_{j+M}^{\mu}=\hat{S}_{j}^{\mu}$ is imposed and we change the basis $(\hat{S}_{j}^{x},\hat{S}_{j}^{y},\hat{S}_{j}^{z})\to(\hat{S}_{j}^{z},\hat{S}_{j}^{x},\hat{S}_{j}^{y})$ to simplify the calculations. Here, we consider only the case of even $M$ for simplicity. A direct calculation shows that the Hamiltonian $\hat{H}_{\rm DH}^{\rm PBC}$ and the following operator and state satisfy the RSGA relations (8), (9), and (10):
$\hat{Q}^{\dagger}$ $\equiv\sum_{j=1}^{M}\hat{P}_{j-1}\hat{S}_{j}^{+}\hat{P}_{j+1},$ (14)
$\left|\psi_{0}\right\rangle$ $\equiv\left|\downarrow_{1}\downarrow_{2}\ldots\downarrow_{M}\right\rangle\equiv\left|\Downarrow\right\rangle,$ (15)
where $\hat{P}_{j}\equiv 1/2-\hat{S}_{j}^{z}$ and $\hat{S}_{j}^{\pm}\equiv\hat{S}_{j}^{x}\pm i\hat{S}_{j}^{y}$. The QMBS states are given by
$\left|S_{n}\right\rangle$ $\propto(\hat{Q}^{\dagger})^{n}\left|\Downarrow\right\rangle,\;n=1,2,\ldots,M/2,$ (16)
which satisfy
$\hat{H}_{\rm DH}^{\rm PBC}\left|S_{n}\right\rangle$ $=\left(\frac{1}{2}M-n\right)h^{x}\left|S_{n}\right\rangle.$ (17)
The operator $\hat{Q}^{\dagger}$ generates the transition $\left|\downarrow_{j-1}\downarrow_{j}\downarrow_{j+1}\right\rangle\to\left|\downarrow_{j-1}\uparrow_{j}\downarrow_{j+1}\right\rangle$. From this fact, we can write down the QMBS states as
$\left|S_{n}\right\rangle$ $=\frac{1}{n!\sqrt{N(M,n)}}(\hat{Q}^{\dagger})^{n}\left|\Downarrow\right\rangle$
$=\frac{1}{\sqrt{N(M,n)}}\sideset{}{{}^{\prime}}{\sum}_{\bm{m}}\left|\bm{m}\right\rangle,$ (18)
$N(M,n)$ $\equiv\frac{M}{M-n}\binom{M-n}{n},$ (19)
where $\left|\bm{m}\right\rangle\equiv\left|m_{1}m_{2}\ldots m_{M}\right\rangle$ represents the basis state, $m_{i}=\uparrow\text{ or }\downarrow$, ${\sideset{}{{}^{\prime}}{\sum}_{\bm{m}}}$ represents the summation over all possible configurations of $\bm{m}$ where $n$ up spins are not adjacent to each other, and $N(M,n)$ is the normalization constant. We also show that the DH model has another series of QMBS states. We define
$\hat{Q}^{\prime\dagger}$ $\equiv\sum_{j=1}^{M}\hat{P}^{\prime}_{j-1}\hat{S}_{j}^{-}\hat{P}^{\prime}_{j+1},$ (20)
$\hat{P}_{j}^{\prime}$ $\equiv\frac{1}{2}+\hat{S}_{j}^{z},$ (21)
$\left|\psi_{0}\right\rangle$ $\equiv\left|\uparrow_{1}\uparrow_{2}\ldots\uparrow_{M}\right\rangle\equiv\left|\Uparrow\right\rangle.$ (22)
Similarly, we can show that the above operators and the states satisfy the RSGA relations. The QMBS states are given by
$\left|S_{n}^{\prime}\right\rangle$ $=\frac{1}{n!\sqrt{N(M,n)}}(\hat{Q}^{\prime\dagger})^{n}\left|\Uparrow\right\rangle=\frac{1}{\sqrt{N(M,n)}}\sideset{}{{}^{\prime\prime}}{\sum}_{\bm{m}}\left|\bm{m}\right\rangle,$ (23)
where ${\sideset{}{{}^{\prime\prime}}{\sum}_{\bm{m}}}$ represents the summation over all possible configurations of $\bm{m}$ where $n$ down spins are not adjacent to each other. The state $\left|S_{n}^{\prime}\right\rangle$ satisfies
$\hat{H}_{\rm DH}^{\rm PBC}\left|S_{n}^{\prime}\right\rangle=\left(-\frac{M}{2}+n\right)h^{x}\left|S_{n}^{\prime}\right\rangle.$ (24)
These two towers of QMBS states are related to each other as $\hat{C}_{x}\left|S_{n}^{\prime}\right\rangle=\left|S_{n}\right\rangle$ through the unitary operator $\hat{C}_{x}\equiv\prod_{j=1}^{M}(2\hat{S}_{j}^{x})$. For $n=M/2$, we can easily show $\left|S_{M/2}\right\rangle=|S_{M/2}^{\prime}\rangle=(\left|\uparrow\downarrow\ldots\uparrow\downarrow\right\rangle+\left|\downarrow\uparrow\ldots\downarrow\uparrow\right\rangle)/\sqrt{2}$. We note that the QMBS states (16) and (23) are almost the same as those with a three-spin interaction found in Refs. Iadecola2020 ; Mark2020 ; Omiya2023 , while our Hamiltonian (13) is different from theirs. Compared to these studies, the DH model has advantages regarding experimental feasibility footnote_DH . We also find that the pyramid scar states. See the Supplementary Material (SM) for a detailed calculation supple .
We now discuss the symmetry of the DH model. We can show that the following operators commute with the Hamiltonian (13):
$\hat{N}_{\rm sol}^{\rm PBC}$ $\equiv\sum_{j=1}^{M}\left(\frac{1}{4}-\hat{S}_{j}^{z}\hat{S}_{j+1}^{z}\right),$ (25)
$\hat{C}$$\equiv\hat{\mathcal{I}}\hat{C}_{z},$ (26)
where $\hat{\mathcal{I}}$ is the space-inversion operator defined by ^$\hat{\mathcal{I}}\hat{S}_{j}^{\mu}\hat{\mathcal{I}}=\hat{S}_{M-j+1}^{\mu}$, and $\hat{C}_{z}\equiv\prod_{j=1}^{M}(2\hat{S}_{j}^{z})$. The operator $\hat{N}_{\rm sol}^{\rm PBC}$ counts the number of solitons Kodama_thesis ; Kodama2023 as discussed in Sec. III. For example, we obtain
[…]
By direct calculations, we have
$\hat{N}_{\rm sol}^{\rm PBC}\left|S_{n}\right\rangle$ $=n\left|S_{n}\right\rangle,\quad\hat{C}\left|S_{n}\right\rangle=(-1)^{n}\left|S_{n}\right\rangle,$ (29)
where we used the relations $[\hat{N}_{\rm sol}^{\rm PBC},\hat{Q}^{\dagger}]=\hat{Q}^{\dagger}$ and $\hat{C}\hat{Q}^{\dagger}\hat{C}=-\hat{Q}^{\dagger}$. Therefore, the QMBS state $\left|S_{n}\right\rangle$ is in the symmetry sector specified by $(N_{\rm sol}^{\rm PBC},C)=(n,(-1)^{n})$.
V.2 Open boundary case
In this subsection, we consider QMBS states under the open boundary condition, where the Hamiltonian reads
$\hat{H}_{\rm DH}^{\rm OBC}$
$\equiv D\sum_{j=1}^{M-1}(\hat{S}_{j}^{z}\hat{S}_{j+1}^{x}-\hat{S}_{j}^{x}\hat{S}_{j+1}^{z})-h^{x}\sum_{j=1}^{M}\hat{S}_{j}^{z}-\frac{D}{2}(\hat{S}_{1}^{x}-\hat{S}_{M}^{x})$ (30a)
$=D\sum_{j=1}^{M}\hat{S}_{j}^{x}(\hat{S}_{j-1}^{z}-\hat{S}_{j+1}^{z})-h^{x}\sum_{j=1}^{M}\hat{S}_{j}^{z}.$ (30b)
Here, we defined $\hat{S}_{0}^{z}=\hat{S}_{M+1}^{z}=-1/2$. The last term in Eq. (30a) is necessary to define the soliton number operator that commutes with Eq. (30a). In Eq. (30b), we add the auxiliary sites $j=0$, $M+1$ and restrict the states to those with the down spins at the two sites. This restriction is equivalent to applying a downward magnetic field with infinite strength to the spins at the auxiliary sites.
Because the Hamiltonian is formally the same as in the periodic case, the QMBS states exist for Eq. (30b) with a slight modification. We modify the definition of $\hat{Q}^{\dagger}$ and $\hat{P}_{j}$ as
$\hat{Q}^{\dagger}$ $\equiv\hat{S}_{1}^{+}\hat{P}_{2}+\sum_{j=2}^{M-1}\hat{P}_{j-1}\hat{S}_{j}^{+}\hat{P}_{j+1}+\hat{P}_{M-1}\hat{S}_{M}^{+}$
$=\sum_{j=1}^{M}\hat{P}_{j-1}\hat{S}_{j}^{+}\hat{P}_{j+1},$ (31)
$\hat{P}_{j}$ $\equiv\frac{1}{2}-\hat{S}_{j}^{z},\quad\hat{P}_{0}\equiv 1,\quad\hat{P}_{M+1}\equiv 1.$ (32)
[…]
$\left|S_{n}\right\rangle$ $\propto(\hat{Q}^{\dagger})^{n}\left|\Downarrow\right\rangle,$ (33)
which satisfies
$\hat{H}_{\rm DH}^{\rm OBC}\left|S_{n}\right\rangle$ $=\left(\frac{1}{2}M-n\right)h^{x}\left|S_{n}\right\rangle.$ (34)
The explicit expression of the QMBS state is given by
$\left|S_{n}\right\rangle$ $=\frac{1}{\sqrt{\mathcal{N}(M,n)}}\sideset{}{{}^{\prime}}{\sum}_{\bm{m}}\left|\bm{m}\right\rangle,$ (35)
$\mathcal{N}(M,n)$ $\equiv\binom{M-n+1}{n},$ (36)
where the meaning of the summation is the same as the case of the periodic boundary condition. We can also show that the Hamiltonian commutes with $\hat{C}$ [the definition is the same as the periodic boundary case, see Eq. (18).] and modified soliton number operator:
$\hat{N}_{\rm sol}^{\rm OBC}$
$\equiv\sum_{j=1}^{M-1}\left(\frac{1}{4}-\hat{S}_{j}^{z}\hat{S}_{j+1}^{z}\right)+\left(\frac{1}{4}+\frac{1}{2}\hat{S}_{1}^{z}\right)+\left(\frac{1}{4}+\frac{1}{2}\hat{S}_{M}^{z}\right)$
$=\sum_{j=0}^{M}\left(\frac{1}{4}-\hat{S}_{j}^{z}\hat{S}_{j+1}^{z}\right).$ (37)
This operator counts the number of solitons in the extended Hilbert space spanned by $\left|\downarrow_{0};\bm{m};\downarrow_{M+1}\right\rangle$. The QMBS states are in the symmetry sector specified by $(N_{\rm sol}^{\rm OBC},C)=(n,(-1)^{n})$.
V.3 Level spacing statistics
Figure 6: $r$-value as a function of the magnetic field. The green solid and black dashed lines represent the $r$-value of the Wigner-Dyson (Gaussian orthogonal ensemble) and Poisson distribution, respectively. (a) DH model without the edge magnetic field for $M=16$. (b) DH model with the edge magnetic field for $M=18$ in $N_{\rm sol}^{\rm OBC}=5$ sector.
Figure 7: $r$-value as a function of the magnetic field in the low-field regime. The green solid and black dashed lines represent the $r$-value of the Wigner-Dyson (Gaussian orthogonal ensemble) and Poisson distribution, respectively. (a) DH model without the edge magnetic field for $M=16$. (b) DH model with the edge magnetic field for $M=18$ in $N_{\rm sol}^{\rm OBC}=5$ sector.
[…]
Figure 6 (a) shows the results of the magnetic-field dependence of the $r$-value of the DH model without an edge magnetic field. In this case, $\hat{C}$ is conserved, but $\hat{N}_{\rm sol}^{\rm OBC}$ is not. The $r$-value is consistent with the Wigner-Dyson distribution (Gaussian orthogonal ensemble) around $h^{x}\lesssim D$ ($r\simeq 0.53$). In the high-field regime, the $r$-value approaches the Poisson value ($r\simeq 0.38$), due to the emergent conservation law of $\hat{S}_{\rm tot}^{z}$ in the high-field regime. We also plot the $r$-value of the DH model with the edge magnetic field in Fig. 6 (b), where we calculate the largest symmetry sector. For $M=18$, e.g,, $N_{\rm sol}^{\rm OBC}=5$ is the largest.
Figures 7 (a) and (b) represent the $r$-value in the low-field regime. We can see the deviation of the $r$-value from the Wigner-Dyson value around $h^{x}\sim 0.01D$. This behavior results from the integrability at $h^{x}=0$; the DH model is equivalent to the XY model at this point in the absence of the edge magnetic field. The integrability of the XY model with the edge magnetic field terms was discussed in Refs. Bilstein1999 ; Bilstein2000 ; Bilstein2000_2 . Regardless of the presence/absence of the edge magnetic field, the $r$-value does not approach the Poisson value. We may attribute this discrepancy to the additional emergent symmetry at $h^{x}=0$ discussed in Refs. Bilstein1999 ; Bilstein2000 ; Bilstein2000_2 .
V.4 Entanglement entropy and off-diagonal long-range order
In this subsection, we discuss the half-chain entanglement entropy of the system under the open boundary condition. Figure 8 (a) shows the half-chain von Neumann EE as a function of the eigenenergy of the system. We find a QMBS state with a high energy eigenstate and low EE. One can derive the analytical expression of the EE of the QMBS state (see Appendix E for details). We show the size dependence of the EE in Fig. 8 (b). The EE obeys the sub-volume law $\sim\ln M$.
[…]
Then, we discuss the off-diagonal long-range order (ODLRO) Yang1962 of the QMBS states. In our case, the order parameter is given by $\hat{Q}^{\dagger}$ [see Eq. (31)]. We can detect ODLRO by the correlation function $\left\langle S_{n}\right|\hat{Q}\hat{Q}^{\dagger}\left|S_{n}\right\rangle$ Tasaki_book . Using the relation
$\hat{Q}^{\dagger}\left|S_{n}\right\rangle$ $=\sqrt{\frac{(n+1)(M-2n+1)(M-2n)}{M-n+1}}\left|S_{n+1}\right\rangle,$ (38)
we can obtain
$\frac{1}{M^{2}}\left\langle S_{n}\right|\hat{Q}\hat{Q}^{\dagger}\left|S_{n}\right\rangle$ $=\frac{(n+1)(M-2n+1)(M-2n)}{M^{2}(M-n+1)}.$ (39)
Here, we fix $n/M=O(1)$. In the thermodynamic limit, Eq. (39) becomes
$\lim_{M\to\infty,n/M;\text{fixed}}\frac{1}{M^{2}}\left\langle S_{n}\right|\hat{Q}\hat{Q}^{\dagger}\left|S_{n}\right\rangle$
$=\frac{n}{M}\left(1-\frac{n/M}{1-n/M}\right)=O(1).$ (40)
Namely, the QMBS states exhibit ODLRO, which we can interpret as magnon condensation Iadecola2020 . Equation (40) indicates that the DH model violates the ETH because ODLRO is not allowed for ETH-obeying states.
V.5 Experimental implications
[…]
VI Asymptotic quantum many-body scar states of the DH model
Recently, Gotta et al. proposed a novel class of nonergodic quantum states, dubbed the AQMBS states Gotta2023 , which are characterized by the following properties: (i) orthogonality to any exact QMBS state, (ii) small entanglement, and (iii) the energy variance that vanishes in the thermodynamics limit. Due to the property (iii), the relaxation time diverges in the thermodynamics limit. According to Ref. Gotta2023 , the AQMBS states can be constructed by deforming the QMBS state. Motivated by Ref. Gotta2023 , we construct the AQMBS states in the DH model. While they can be constructed for both periodic and open boundary conditions, we focus on the latter in the following. See the SM for details supple .
The AQMBS states in the DH model are given by
$\left|AS_{n}\right\rangle$ $\propto\hat{A}^{\dagger}\left|S_{n-1}\right\rangle,\quad n=1,2,\ldots,M/2,$ (41)
$\hat{A}^{\dagger}$ $\equiv\sum_{j=1}^{M}f_{j}\hat{P}_{j-1}\hat{S}_{j}^{+}\hat{P}_{j+1},$ (42)
where $f_{j}$ is a complex function that satisfies the condition $f_{j}=-f_{M+1-j}$. This condition ensures $\left\langle AS_{n}|S_{m}\right\rangle=0$ because $\left|AS_{n}\right\rangle$ and $\left|S_{n}\right\rangle$ have different parity under spatial inversion. Here, we choose $f_{j}=\cos[\pi j/(M+1)]$, but other choices are possible (see the SM supple ).
To verify that the AQMBS states have a small amount of entanglement, we study the half-chain EE of Eq. (41), which we can calculate from the matrix product state (MPS) representation Schollwock2011 ; Paeckel2019 of Eq. (41). Our analytical result shows that the bond dimension of the state (41) is at most $\chi=n(3n+1)$, which is of the order of $O(M^{2})$ because $n$ is at most $O(M)$. The half-chain EE is bounded by $\ln\chi\sim\ln M$. Therefore, the state (41) obeys a sub-volume law scaling.
Figure 10 shows numerical results of the half-chain EE of the states $\left|AS_{n}\right\rangle$ for several $n$ and system sizes. We see that the system size dependence of the half-chain EE is weak and their values are small compared to the Page value Page1993 , which is given by $S_{\rm Page}=(M/2)\ln(2)-1/2\simeq 86$ for $M=250$.
Figure 10: System size dependence of the half-chain EE for the states $\left|AS_{n}\right\rangle$ for several $n$.
Next, we estimate the energy variance of the state (41), which we can write as
$\Delta E^{2}$ $\equiv\left\langle AS_{n}\right|(\hat{H}_{\rm DH}^{\rm OBC})^{2}\left|AS_{n}\right\rangle-(\left\langle AS_{n}\right|\hat{H}_{\rm DH}^{\rm OBC}\left|AS_{n}\right\rangle)^{2}$
$=\frac{\left\langle S_{n-1}\right|[\hat{A},\hat{H}_{\rm DM}^{\rm OBC}][\hat{H}_{\rm DM}^{\rm OBC},\hat{A}^{\dagger}]\left|S_{n-1}\right\rangle}{\left\langle S_{n-1}\right|\hat{A}\hat{A}^{\dagger}\left|S_{n-1}\right\rangle},$ (43)
where we defined $\hat{H}_{\rm DM}^{\rm OBC}\equiv D\sum_{j=1}^{M}\hat{S}_{j}^{x}(\hat{S}_{j-1}^{z}-\hat{S}_{j+1}^{z})$ and used the fact that $\hat{H}_{\rm DM}^{\rm OBC}\left|S_{n-1}\right\rangle=0$ and the state (41) is an eigenstate of the Zeeman term. For $n=O(1)$, we can exactly evaluate the energy variance as $\Delta E^{2}=\pi^{2}D^{2}/(4M^{2})+O(M^{-3})$. We show the numerical results of the energy variance in Fig. 11. For other $n$, we numerically verify that the energy variance goes to zero in the thermodynamic limit (see the SM for details supple .).
Figure 11: The energy variance of the states $\left|AS_{n}\right\rangle$ for several $n$ as a function of the system size $M$. “Leading" in the legend means $\Delta E^{2}/D^{2}=\pi^{2}/(4M^{2})$.
Finally, we discuss the dynamics after quench from the state $\left|AS_{n}\right\rangle$. Figures 12 (a), (b), and (c) show the fidelity $|\left\langle AS_{n}\right|e^{-i\hat{H}_{\rm DH}^{\rm OBC}t/\hbar}\left|AS_{n}\right\rangle|^{2}$ for several system sizes. We use the TEBD method to obtain these results. The discrete time step is set to $\Delta t=0.01\hbar/D$, and the bond dimension is taken to be $\chi=200\sim 310$ so that the truncation error is less than $10^{-10}$. At early times, the fidelity can be approximated as $|\left\langle AS_{n}\right|\exp(-i\hat{H}_{\rm DH}^{\rm OBC}t/\hbar)\left|AS_{n}\right\rangle|^{2}\simeq 1-\Delta E^{2}t^{2}/\hbar^{2}$, which implies that the smaller energy variance yields a longer relaxation time. As expected, we see the system size dependence of the fidelity even for $n>1$. We also plot the rescaled fidelity in Figs. 12(d), (e), and (f). We can find a good data collapse by rescaling $t/M$. These results all suggest that the state (41) is an AQMBS state in the DH model.
Figure 12: Time evolution of the fidelity with the initial condition (a) $\left|AS_{1}\right\rangle$, (b) $\left|AS_{5}\right\rangle$, (c) $\left|AS_{10}\right\rangle$. (d), (e), and (f): Rescaled fidelity of (a), (b), and (c), respectively.
VII Summary
In this paper, we proposed a new method for performing quantum simulations of the $S=1/2$ quantum spin model with the DMI using Rydberg atoms. Our approach is based on the two-photon Raman scheme and unitary transformation, which is feasible with current experimental techniques. One advantage of the Rydberg atom quantum simulator is that it realizes quantum spin models with high tunability. In our scheme, we can easily tune the ratio between the magnitude of the exchange interaction and the DMI in a wide range. This tunability allows us to implement the DH model, which has only the DMI and Zeeman energy. Another advantage is that the heating problem is insignificant compared to the SOC case because the typical timescale is much shorter than the neutral atom case due to the strong interaction between the Rydberg atoms.
We studied the ground-state properties and nonequilibrium dynamics of the DH model. We showed that the time-averaged magnetization continuously changes as a function of the magnetic field in the fully quantum case, whereas it is discontinuous in the classical case. This behavior may originate from the macroscopic quantum tunneling due to the strong quantum fluctuations.
We have also investigated the nonergodic properties of the DH model. Using the RSGA formalism, we analytically demonstrated the existence of QMBS states in the DH model with open and periodic boundary conditions. Furthermore, we identified a series of AQMBS states in the DH model. These states are orthogonal to all exact QMBS states, have low entanglement, and exhibit slow relaxation due to their small energy variance.
Our work opens a new avenue for the design of quantum chiral magnets. There are several future directions. One is to extend our scheme to a more complex form of the DMI, which would enable experimental studies of skyrmions in the quantum regime Ochoa2019 . Another direction is an extension to higher spin systems, allowing us to explore the spin parity effect in the DH model.
Acknowledgements.
We thank D. Yamamoto, T. Nikuni, and S. Moudgalya for their helpful comments. This work was supported by JSPS KAKENHI Grants No. JP20K14389 (M.K.), No. JP22H05268 (M.K.), No. JP22H05267 (T.T.), No. JP21H05191 (H.K.), No. JP18K03445 (H.K.), No. JP23H01093 (H.K.), No. JP20K03855 (Y.K.), and No. JP21H01032 (Y.K.). This research was also supported by Joint Research by the Institute for Molecular Science (IMS program No. 23IMS1101) (Y.K.)
Appendix A Derivation of the XXZ Hamiltonian
In this appendix, we discuss the interaction between Rydberg states $|n_{1}S_{1/2},m_{J}=+1/2\rangle\equiv\left|\downarrow\right\rangle$ and $|n_{2}S_{1/2},m_{J}=+1/2\rangle\equiv\left|\uparrow\right\rangle$ with energy $E_{n_{1}S_{1/2}}$ and $E_{n_{2}S_{1/2}}$. Since these two states have the same parity, they are not coupled directly via the dipole-dipole interaction. If the atomic distance is large enough, the dipole-dipole interaction can be treated as a perturbation, and the leading term is given by the second-order term Bijnen_thesis ; Whitlock2017 . For simplicity, we consider only the two Rydberg atoms separated by the distance $R$. The Hamiltonian is given by Weber2017
$\hat{H}_{\rm two}$ $=\hat{H}_{\rm Rydberg}+\hat{V}_{\rm dip},$ (44)
$\hat{H}_{\rm Rydberg}$ $=\sum_{n,l,J,m_{J}}E_{n,l,J,m_{J}}(\left|n_{,}l,J,m_{J}\right\rangle\left\langle n,l,J,m_{J}\right|\otimes\hat{1}_{2})$
$\quad+\sum_{n^{\prime},l^{\prime},J^{\prime},m_{J}^{\prime}}E_{n^{\prime},l^{\prime},J^{\prime},m_{J}^{\prime}}(\hat{1}_{1}\otimes\left|n^{\prime},l^{\prime},J^{\prime},m_{J}^{\prime}\right\rangle\left\langle n^{\prime},l^{\prime},J^{\prime},m_{J}^{\prime}\right|),$ (45)
$\hat{V}_{\rm dip}$ $=\frac{1}{4\pi\epsilon_{0}R^{3}}\left[\hat{\bm{d}}_{1}\cdot\hat{\bm{d}}_{2}-3(\hat{\bm{d}_{1}}\cdot\tilde{\bm{R}})(\hat{\bm{d}}_{2}\cdot\tilde{\bm{R}})\right],$ (46)
[…]
$\mathcal{E}_{1}$ $\equiv 2E_{n_{2}S_{1/2}},\quad\mathcal{E}_{2}\equiv\mathcal{E}_{3}\equiv E_{n_{2}S_{1/2}}+E_{n_{1}S_{1/2}},\quad\mathcal{E}_{4}\equiv 2E_{n_{1}S_{1/2}}.$ (47)
Using the standard second-order perturbation theory, we obtain the effective Hamiltonian:
$\hat{H}_{\rm two}^{(2)}$ $=\sum_{j=1}^{4}\left(\mathcal{E}_{j}\left|j\right\rangle\left\langle j\right|\right)+\left[J_{11}\left|1\right\rangle\left\langle 1\right|+J_{22}\left|2\right\rangle\left\langle 2\right|+J_{33}\left|3\right\rangle\left\langle 3\right|+J_{44}\left|4\right\rangle\left\langle 4\right|+\frac{J}{2}(\left|2\right\rangle\left\langle 3\right|+\left|3\right\rangle\left\langle 2\right|)\right],$ (48)
$J_{11}$ $\equiv-\sum_{\begin{subarray}{c}n,l,J,m_{J},n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}\end{subarray}}\frac{\left\langle 1\right|\hat{V}_{\rm dip}\left|n,l,J,m_{J},n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}\right\rangle\left\langle n,l,J,m_{J},n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}\right|\hat{V}_{\rm dip}\left|1\right\rangle}{E_{n,l,J,m_{J}}+E_{n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}}-\mathcal{E}_{1}}\equiv\frac{C_{6}(n_{2},n_{2})}{R^{6}},$ (49)
$J_{22}$ $=J_{33}\equiv-\sum_{\begin{subarray}{c}n,l,J,m_{J},n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}\end{subarray}}\frac{\left\langle 2\right|\hat{V}_{\rm dip}\left|n,l,J,m_{J},n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}\right\rangle\left\langle n,l,J,m_{J},n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}\right|\hat{V}_{\rm dip}\left|2\right\rangle}{E_{n,l,J,m_{J}}+E_{n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}}-\mathcal{E}_{2}}\equiv\frac{C_{6}(n_{1},n_{2})}{R^{6}},$ (50)
$J_{44}$ $\equiv-\sum_{\begin{subarray}{c}n,l,J,m_{J},n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}\end{subarray}}\frac{\left\langle 4\right|\hat{V}_{\rm dip}\left|n,l,J,m_{J},n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}\right\rangle\left\langle n,l,J,m_{J},n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}\right|\hat{V}_{\rm dip}\left|4\right\rangle}{E_{n,l,J,m_{J}}+E_{n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}}-\mathcal{E}_{4}}\equiv\frac{C_{6}(n_{1},n_{1})}{R^{6}},$ (51)
$\frac{J}{2}$ $\equiv-\sum_{\begin{subarray}{c}n,l,J,m_{J},n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}\end{subarray}}\frac{\left\langle 2\right|\hat{V}_{\rm dip}\left|n,l,J,m_{J},n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}\right\rangle\left\langle n,l,J,m_{J},n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}\right|\hat{V}_{\rm dip}\left|3\right\rangle}{E_{n,l,J,m_{J}}+E_{n^{\prime},l^{\prime},J^{\prime},m^{\prime}_{J^{\prime}}}-\mathcal{E}_{2}}\equiv\frac{\tilde{C}_{6}(n_{1},n_{2})}{R^{6}}.$ (52)
Because the state $\left|j\right\rangle(j=1,2,3,4)$ is an $S$ state, the only nonvanishing intermediate states are $P$ states due to the selection rule. We can write down the effective Hamiltonian in the spin language. The spin operators can be defined as
$\hat{S}_{i}^{+}$ $\equiv\left|\uparrow_{i}\right\rangle\left\langle\downarrow_{i}\right|,$ (53)
$\hat{S}_{i}^{-}$ $\equiv\left|\downarrow_{i}\right\rangle\left\langle\uparrow_{i}\right|,$ (54)
$\hat{S}_{i}^{z}$ $\equiv\frac{1}{2}(\left|\uparrow_{i}\right\rangle\left\langle\uparrow_{i}\right|-\left|\downarrow_{i}\right\rangle\left\langle\downarrow_{i}\right|),$ (55)
$\hat{S}_{i}^{x}$ $\equiv(\hat{S}_{i}^{+}+\hat{S}_{i}^{-})/2,\quad\hat{S}_{i}^{y}\equiv(\hat{S}_{i}^{+}-\hat{S}_{i}^{-})/(2i).$ (56)
From the definitions of the spin operators and Eq. (48), we obtain the XXZ type Hamiltonian
$\hat{H}_{\rm two}^{(2)}$ $=\left[E_{n_{2}S_{1/2}}-E_{n_{1}S_{1/2}}+\frac{J_{11}-J_{44}}{2}\right](\hat{S}_{1}^{z}+\hat{S}_{2}^{z})+J(\hat{S}_{1}^{x}\hat{S}_{2}^{x}+\hat{S}_{1}^{y}\hat{S}_{2}^{y}+\delta\hat{S}_{1}^{z}\hat{S}_{2}^{z})+\frac{1}{4}(J_{11}+J_{44}+2J_{22}),$ (57)
$\delta$ $\equiv\frac{J_{11}+J_{44}-2J_{22}}{J}=\frac{C_{6}(n_{1},n_{1})+C_{6}(n_{2},n_{2})-2C_{6}(n_{1},n_{2})}{\tilde{C}_{6}(n_{1},n_{2})}.$ (58)
We then extend the above results to the case of $M$ atoms. We assume that $M$ Rydberg atoms are arranged in a one-dimensional open chain. In this case, the effective Hamiltonian is given by
$\hat{H}_{\rm multi}^{(2)}$ =$\frac{1}{2}\sum_{j,k,j\not=k}J_{jk}(\hat{S}_{j}^{x}\hat{S}_{k}^{x}+\hat{S}_{j}^{y}\hat{S}_{k}^{y}+\delta\hat{S}_{j}^{z}\hat{S}_{k}^{z})+\sum_{j=1}^{M}h_{j}^{z}\hat{S}_{j}^{z},$ (59)
where $R_{jk}\equiv d|j-k|$, $J_{jk}\equiv\tilde{C}_{6}(n_{1},n_{2})/R_{jk}^{6}$, and $h_{j}^{z}\equiv(E_{n_{2}S_{1/2}}-E_{n_{1}S_{1/2}})+\left\{[C_{6}(n_{2},n_{2})-C_{6}(n_{1},n_{1})]/2\right\}\sum_{k\not=j}1/R_{jk}^{6}$. In this work, we neglect the second term of $h_{j}^{z}$ because this term is small compared with the first term in typical experimental situations.
Appendix B Derivation of the rotating transverse field term
In this appendix, we consider the single Rydberg atom at the position $\bm{R}_{j}\equiv(R_{jx},R_{jy},R_{jz})\equiv dj(\cos\theta,0,\sin\theta)$ (see Fig. 1). We irradiate two linearly polarized lasers with detuning $\hbar\Delta_{1}=\hbar\omega_{1}-(E_{n_{1}S_{1/2}}-E_{P})$ and $\hbar\Delta_{2}=\hbar\omega_{2}-(E_{n_{2}S_{1/2}}-E_{P})$, where $E_{P}$ is the energy of the intermediate state $\left|6P_{3/2}\right\rangle$. The atom-light interaction Hamiltonian under the rotating wave approximation can be written as
$\hat{H}_{\rm AL}(t)$ $=E_{n_{1}S_{1/2}}|n_{1}S_{1/2}\rangle\langle n_{1}S_{1/2}|+E_{n_{2}S_{1/2}}|n_{2}S_{1/2}\rangle\langle n_{2}S_{1/2}|+E_{P}|6P_{3/2}\rangle\langle 6P_{3/2}|$
$\quad+\frac{\hbar\Omega_{1}}{2}\left[e^{+i(k_{1}R_{jx}-\omega_{1}t)}|n_{1}S_{1/2}\rangle\langle 6P_{3/2}|+e^{-i(k_{1}R_{jx}-\omega_{1}t)}|6P_{3/2}\rangle\langle n_{1}S_{1/2}|\right]$
$\quad+\frac{\hbar\Omega_{2}}{2}\left[e^{+i(k_{2}R_{jy}-\omega_{2}t)}|n_{2}S_{1/2}\rangle\langle 6P_{3/2}|+e^{-i(k_{2}R_{jy}-\omega_{2}t)}|6P_{3/2}\rangle\langle n_{2}S_{1/2}|\right].$ (60)
Here, we omit the values of $m_{J}$. To transform into the rotating frame of the laser field, we define $\hat{H}_{0}$ as
$\hat{H}_{0}$ $=\left(E_{n_{1}S_{1/2}}-\hbar\tilde{\omega}\right)|n_{1}S_{1/2}\rangle\langle n_{1}S_{1/2}|+(E_{n_{1}S_{1/2}}-\hbar\omega_{1}-\hbar\tilde{\omega})|6P_{3/2}\rangle\langle 6P_{3/2}|$
$\quad+(E_{n_{1}S_{1/2}}-\hbar\omega_{1}+\hbar\omega_{2}-\hbar\tilde{\omega})|n_{2}S_{1/2}\rangle\langle n_{2}S_{1/2}|,$ (61)
$\hbar\tilde{\omega}$ $\equiv\frac{\hbar}{2}(-\Delta_{1}+\Delta_{2}).$ (62)
The Hamiltonian in the rotating frame of the laser field is given by
$\hat{H}_{\rm AL;rot}$ $\equiv e^{i\hat{H}_{0}t/\hbar}[\hat{H}_{\rm AL}(t)-\hat{H}_{0}]e^{-i\hat{H}_{0}t/\hbar}$
$=-\frac{\hbar}{2}(\Delta_{1}-\Delta_{2})|n_{1}S_{1/2}\rangle\langle n_{1}S_{1/2}|+\frac{\hbar}{2}(\Delta_{1}+\Delta_{2})|6P_{3/2}\rangle\langle 6P_{3/2}|+\frac{\hbar}{2}(\Delta_{1}-\Delta_{2})|n_{2}S_{1/2}\rangle\langle n_{2}S_{1/2}|$
$\quad+\frac{\hbar\Omega_{1}}{2}\left(e^{+ik_{1}R_{jx}}|n_{1}S_{1/2}\rangle\langle 6P_{3/2}|+e^{-ik_{1}R_{jx}}|6P_{3/2}\rangle\langle n_{1}S_{1/2}|\right)$
$\quad+\frac{\hbar\Omega_{2}}{2}\left(e^{+ik_{2}R_{jy}}|n_{2}S_{1/2}\rangle\langle 6P_{3/2}|+e^{-ik_{2}R_{jy}}|6P_{3/2}\rangle\langle n_{2}S_{1/2}|\right).$ (63)
From now on, we assume that $|\Delta_{1}-\Delta_{2}|\ll|\Delta_{1}|,|\Delta_{2}|$ and use the standard adiabatic elimination scheme Steck ; Paulisch2014 or Schrieffer-Wolff transformation CohenTannoudjibook ; Bravyi2011 . We then obtain the effective Hamiltonian for $j$th atom:
$\hat{h}_{j}$ $=-\hbar\left[\frac{\Delta_{1}-\Delta_{2}}{2}+\frac{\Omega_{1}^{2}}{4(\Delta_{1}+\Delta_{2})/2}\right]|n_{1}S_{1/2}\rangle\langle n_{1}S_{1/2}|+\hbar\left[\frac{\Delta_{1}-\Delta_{2}}{2}-\frac{\Omega_{2}^{2}}{4(\Delta_{1}+\Delta_{2})/2}\right]|n_{2}S_{1/2}\rangle\langle n_{2}S_{1/2}|$
$\quad-\hbar\frac{\Omega_{1}\Omega_{2}}{4(\Delta_{1}+\Delta_{2})/2}e^{+ik_{1}R_{jx}}|n_{1}S_{1/2}\rangle\langle n_{2}S_{1/2}|-\hbar\frac{\Omega_{1}\Omega_{2}}{4(\Delta_{1}+\Delta_{2})/2}e^{-ik_{1}R_{jx}}|n_{2}S_{1/2}\rangle\langle n_{1}S_{1/2}|,$ (64)
where we used $R_{jy}=0$. In the spin language, the effective Hamiltonian becomes
$\hat{h}_{j}$ $=-\hbar\tilde{\Delta}\hat{S}_{j}^{z}-\hbar\Omega_{\rm eff}[\cos(k_{1}dj\cos\theta)\hat{S}_{j}^{x}+\sin(k_{1}dj\cos\theta)\hat{S}_{j}^{y}]-\frac{\hbar(\Omega_{1}^{2}+\Omega_{2}^{2})}{8\Delta},$ (65)
where $\tilde{\Delta}=-[\Delta_{1}-\Delta_{2}+(\Omega_{1}^{2}-\Omega_{2}^{2})/(4\Delta)]$, $\Delta=(\Delta_{1}+\Delta_{2})/2$, and $\Omega_{\rm eff}=\Omega_{1}\Omega_{2}/(2\Delta)$. In the main text, we omit the constant term in Eq. (65).
Appendix C Relation between the spin-laboratory and spin-rotating frames
In this appendix, we discuss the relation between the spin-laboratory and spin-rotating frames. The DMI results from the unitary transformation
$\hat{U}_{\text{s-rot}}$ $=\prod_{j=1}^{M}e^{-i\hat{S}_{j}^{z}qj},\quad q=k_{1}d\cos\theta,$ (66)
where $k_{1}$ is the wave number of laser 1, $d$ is the lattice constant, and $\theta$ is the angle between the chain and laser 1 (see Fig. 1). Under this transformation, the spin operators are transformed to
$\hat{U}_{\text{s-rot}}^{\dagger}\hat{S}_{j}^{x}\hat{U}_{\text{s-rot}}$ $=\hat{S}_{j}^{x}\cos(qj)-\hat{S}_{j}^{y}\sin(qj),$ (67)
$\hat{U}_{\text{s-rot}}^{\dagger}\hat{S}_{j}^{y}\hat{U}_{\text{s-rot}}$ $=\hat{S}_{j}^{x}\sin(qj)+\hat{S}_{j}^{y}\cos(qj),$ (68)
$\hat{U}_{\text{s-rot}}^{\dagger}\hat{S}_{j}^{z}\hat{U}_{\text{s-rot}}$ $=\hat{S}_{j}^{z}.$ (69)
From these expressions, we obtain the Hamiltonian in the spin-rotating frame: $\hat{H}_{\text{s-rot}}=\hat{U}^{\dagger}_{\text{s-rot}}\hat{H}_{\text{s-lab}}\hat{U}_{\text{s-rot}}$. A state in the spin-laboratory frame $\left|\psi_{\text{s-lab}}\right\rangle$ is transformed to $\left|\psi_{\text{s-rot}}\right\rangle\equiv\hat{U}^{\dagger}_{\text{s-rot}}\left|\psi_{\text{s-lab}}\right\rangle$.
Then, we consider the expectation values of the spin operators in the spin-laboratory and spin-rotating frames. The expectation value in the laboratory frame is defined by
$\langle\hat{\bm{S}}_{j}\rangle_{\text{s-lab}}$ $\equiv\left\langle\psi_{\text{s-lab}}\right|\hat{\bm{S}}_{j}\left|\psi_{\text{s-lab}}\right\rangle.$ (70)
The expectation values in the spin-rotating frame are defined by
$\langle\hat{\bm{S}}_{j}\rangle_{\text{s-rot}}\equiv\left\langle\psi_{\text{s-rot}}\right|\hat{\bm{S}}_{j}\left|\psi_{\text{s-rot}}\right\rangle.$ (71)
Using the relation $\left|\psi_{\text{s-rot}}\right\rangle=\hat{U}^{\dagger}_{\text{s-rot}}\left|\psi_{\text{s-lab}}\right\rangle$, we obtain $\langle\bm{S}_{j}\rangle_{\text{s-rot}}=\left\langle\psi_{\text{s-lab}}\right|\hat{U}\hat{\bm{S}}_{j}\hat{U}^{\dagger}\left|\psi_{\text{s-lab}}\right\rangle$. Therefore, the relation of the spin expectation value between the spin-laboratory and spin-rotating frames becomes
$\langle\hat{S}_{j}^{x}\rangle_{\text{s-rot}}$ $=\langle\hat{S}_{j}^{x}\rangle_{\text{s-lab}}\cos(qj)+\langle\hat{S}_{j}^{y}\rangle_{\text{s-lab}}\sin(qj),$ (72)
$\langle\hat{S}_{j}^{y}\rangle_{\text{s-rot}}$ $=\langle\hat{S}_{j}^{x}\rangle_{\text{s-lab}}\sin(qj)-\langle\hat{S}_{j}^{y}\rangle_{\text{s-lab}}\cos(qj),$ (73)
$\langle\hat{S}_{j}^{z}\rangle_{\text{s-rot}}$ $=\langle\hat{S}_{j}^{z}\rangle_{\text{s-lab}}.$ (74)
In the actual experiments, the physical quantities are measured in the spin-laboratory frame. From the above relations, we can obtain the physical quantities in the spin-rotating frame.
We can measure the spin expectation values as follows. In typical Rydberg atom experiments with optical tweezers, we can extract $\langle\hat{S}_{j}^{z}\rangle_{\text{s-lab}}$ from a measured recapture probability. To obtain $\langle\hat{S}_{j}^{x}\rangle_{\text{s-lab}}$ and $\langle\hat{S}_{j}^{y}\rangle_{\text{s-lab}}$, the quantum-state tomographic technique is available. That is, we irradiate an appropriate $\pi/2$ pulse to the system before measuring the recapture probability. In our case, we can use a resonant two-photon microwave field as the $\pi/2$ pulse Signoles2021 . We can obtain $\langle\hat{S}_{j}^{x}\rangle_{\text{s-lab}}$ and $\langle\hat{S}_{j}^{y}\rangle_{\text{s-lab}}$ in this way.
For convenience, we summarize the correspondence of typical quantum states and spin expectation values between the spin-laboratory and spin-rotating frames in table 1.
Table 1: List of the correspondence between the spin-laboratory and spin-rotating frames. $\left|\uparrow_{j}\right\rangle$ and $\left|\downarrow_{j}\right\rangle$ represent the eigenstates of $\hat{S}_{j}^{z}$ with eigenvalue $+1/2$ and $-1/2$, respectively.
Spin-laboratory frame Spin-rotating frame
$\left|\psi_{\text{s-lab}}\right\rangle$ $\langle\hat{\bm{S}}_{j}\rangle_{\text{s-lab}}\equiv\left\langle\psi_{\text{s-lab}}\right|\hat{\bm{S}}_{j}\left|\psi_{\text{s-lab}}\right\rangle$ $\left|\psi_{\text{s-rot}}\right\rangle\equiv\hat{U}^{\dagger}_{\text{s-rot}}\left|\psi_{\text{s-lab}}\right\rangle$ $\langle\hat{\bm{S}}_{j}\rangle_{\text{s-rot}}\equiv\left\langle\psi_{\text{s-rot}}\right|\hat{\bm{S}}_{j}\left|\psi_{\text{s-rot}}\right\rangle$
$\left|\uparrow_{j}\right\rangle$ $\dfrac{1}{2}(0,0,+1)$ $\left|\uparrow_{j}\right\rangle$ $\dfrac{1}{2}(0,0,+1)$
$\left|\downarrow_{j}\right\rangle$ $\dfrac{1}{2}(0,0,-1)$ $\left|\downarrow_{j}\right\rangle$ $\dfrac{1}{2}(0,0,-1)$
$\dfrac{1}{\sqrt{2}}(\left|\uparrow_{j}\right\rangle\pm\left|\downarrow_{j}\right\rangle)$ $\dfrac{1}{2}(\pm 1,0,0)$ $\dfrac{1}{\sqrt{2}}(e^{+iqj/2}\left|\uparrow_{j}\right\rangle\pm e^{-iqj/2}\left|\downarrow_{j}\right\rangle)$ $\dfrac{1}{2}(\pm\cos qj,\mp\sin qj,0)$
$\dfrac{1}{\sqrt{2}}(e^{-iqj/2}\left|\uparrow_{j}\right\rangle\pm e^{+iqj/2}\left|\downarrow_{j}\right\rangle)$ $\dfrac{1}{2}(\pm\cos qj,\pm\sin qj,0)$ $\dfrac{1}{\sqrt{2}}(\left|\uparrow_{j}\right\rangle\pm\left|\downarrow_{j}\right\rangle)$ $\dfrac{1}{2}(\pm 1,0,0)$
Appendix D Extension to two-dimensional systems
In this appendix, we discuss an extension of our setup to a two-dimensional system. We consider the setup shown in Fig. 13. The $j$th atom is located at the lattice site $\bm{R}_{j}\equiv d(n_{x,j}\bm{e}_{x}^{\prime}+n_{z,j}\bm{e}_{z}^{\prime})$, where $n_{x,j}$ and $n_{z,j}$ are integers, $\bm{e}_{x}^{\prime}\equiv\cos\theta\bm{e}_{x}+\sin\theta\bm{e}_{z}$, $\bm{e}^{\prime}_{z}\equiv\cos\theta\bm{e}_{z}-\sin\theta\bm{e}_{x}$, and $\bm{e}_{x}$ and $\bm{e}_{z}$ are the unit vectors in the $x$ and $z$ directions, respectively.
Figure 13: Schematic of the setup for the experimental realization of the DMI in a two-dimensional square lattice. The filled black circles represent the position of the Rydberg atoms. The magnetic field is applied along the $z$ axis.
In this case, the total Hamiltonian in the spin-laboratory frame is given by
$\hat{H}_{\text{s-lab}}^{2{\rm D}}$ $=\frac{1}{2}\sum_{\begin{subarray}{c}j,k\\ j\neq k\end{subarray}}J_{jk}(\hat{S}_{j}^{x}\hat{S}_{k}^{x}+\hat{S}_{j}^{y}\hat{S}_{k}^{y}+\delta\hat{S}_{j}^{z}\hat{S}_{k}^{z})-\hbar\Omega_{\rm eff}\sum_{j}\left[\cos(\bm{k}_{1}\cdot\bm{R}_{j})\hat{S}_{j}^{x}+\sin(\bm{k}_{1}\cdot\bm{R}_{j})\hat{S}_{j}^{y}\right]-\hbar\tilde{\Delta}\sum_{j}\hat{S}_{j}^{z},$ (75)
where $J_{jk}\equiv\tilde{C}_{6}(n_{1},n_{2})/|\bm{R}_{j}-\bm{R}_{k}|^{6}$. Using the unitary operator $\hat{U}_{\text{s-rot}}^{\rm 2D}\equiv\prod_{j}e^{-i\hat{S}_{j}^{z}\bm{k}_{1}\cdot\bm{R}_{j}}$, we obtain the Hamiltonian in the spin-rotating frame:
$\hat{H}_{\text{s-rot}}^{\rm 2D}$ $=(\hat{U}^{\rm 2D}_{\text{s-rot}})^{\dagger}\hat{H}_{\text{s-lab}}^{\rm 2D}\hat{U}_{\text{s-rot}}^{\rm 2D}$
$=\frac{1}{2}\sum_{\begin{subarray}{c}j,k\\ j\neq k\end{subarray}}J_{jk}\left\{\cos[\bm{k}_{1}\cdot(\bm{R}_{j}-\bm{R}_{k})](\hat{S}_{j}^{x}\hat{S}_{k}^{x}+\hat{S}_{j}^{y}\hat{S}_{k}^{y})+\delta\hat{S}_{j}^{z}\hat{S}_{k}^{z}+\sin[\bm{k}_{1}\cdot(\bm{R}_{j}-\bm{R}_{k})](\hat{S}_{j}^{x}\hat{S}_{k}^{y}-\hat{S}_{j}^{y}\hat{S}_{k}^{x})\right\}$
$\quad-\hbar\Omega_{\rm eff}\sum_{j}\hat{S}_{j}^{x}-\hbar\tilde{\Delta}\sum_{j}\hat{S}_{j}^{z}.$ (76)
This result shows that the interaction depends on the directions. For example, when $\theta=0$, the DMI appears only for the $x$ direction.
Appendix E Analytical expression of the von Neumann entanglement entropy of the quantum many-body scar states
The QMBS states of the DH model are essentially the same as those found in Ref. Iadecola2020 . In this appendix, we follow their approach and calculate the von Neumann EE of the QMBS states in our model.
We first rewrite Eq. (35) as
$\left|S_{n}\right\rangle=\sum_{\bm{m}_{A},\bm{m}_{B}}M_{\bm{m}_{A},\bm{m}_{B}}^{n}\left|\bm{m}_{A}\right\rangle\left|\bm{m}_{B}\right\rangle,$ (77)
where $\bm{m}_{A(B)}$ denotes a spin configuration in the subsystem $A(B)$, $M^{n}_{\bm{m}_{A},\bm{m}_{B}}$ is a $D_{A}\times D_{B}$ matrix and $D_{A(B)}$ is the dimension of the subsystem $A$ $(B)$. From the explicit expression of the scar state (35), $M^{n}_{\bm{m}_{A},\bm{m}_{B}}$ becomes zero if spin configurations ($\left|\bm{m}\right\rangle=\left|\bm{m}_{A}\right\rangle\left|\bm{m}_{B}\right\rangle$) violate the constraint that up spins are not adjacent to each other. In the following, we consider the half-chain EE, i.e., the sizes of the subsystem $A$ and $B$ are the same. In this case, Eq. (77) reduces to
$\left|S_{n}\right\rangle$ $=\sum_{\bm{m}_{A},\bm{m}_{B}}\sum_{k=0}^{K}M^{n,k}_{\bm{m}_{A},\bm{m}_{B}}\left|\bm{m}_{A}\right\rangle\left|\bm{m}_{B}\right\rangle,$ (78)
where $k$ represents the number of up spins in the subsystem $A$, and $K\equiv\min({n,1+\lfloor M/4\rfloor})$. The element of the matrix $M^{n,k}_{\bm{m}_{A},\bm{m}_{B}}$ takes the value $1/\sqrt{N(M,n)}$ [see Eq. (36)] if $\left|\bm{m}_{A}\right\rangle$ contains $k$-up spins and $\left|\bm{m}_{B}\right\rangle$ contains $(n-k)$-up spins and the constraint of the spin configuration is satisfied, otherwise zero.
To calculate the EE, we consider the matrix element of the reduced density matrix $\rho^{n,k}_{\bm{m}_{A},\bm{m}_{A}^{\prime}}\equiv\sum_{\bm{m}_{B}}M^{n,k}_{\bm{m}_{A},\bm{m}_{B}}(M^{n,k}_{\bm{m}^{\prime}_{A},\bm{m}_{B}})^{\dagger}$. This matrix takes the form
$\rho^{n,k}_{\bm{m}_{A},\bm{m}_{B}}$ $=\frac{1}{\mathcal{N}(M,n)}\begin{bmatrix}l_{1,k}\bm{1}_{D_{1,k}\times D_{1,k}}&l_{2,k}\bm{1}_{D_{1,k}\times D_{2,k}}\\ l_{2,k}\bm{1}_{D_{2,k}\times D_{1,k}}&l_{2,k}\bm{1}_{D_{2,k}\times D_{2,k}}\end{bmatrix},$ (79)
$D_{1,k}$ $\equiv N(M/2-1,k),$ (80)
$D_{2,k}$ $\equiv N(M/2-2,k-1),$ (81)
$l_{1,k}$ $\equiv N(M/2,n-k),$ (82)
$l_{2,k}$ $\equiv N(M/2-1,n-k),$ (83)
where $\bm{1}_{n\times m}$ is an $n\times m$ matrix whose all entries are equal to 1, $D_{1,k}(D_{2,k})$ is the number of possible configurations of the QMBS state in the subsystem $A$ with $m_{M/2}=\downarrow(\uparrow)$. Similarly, $l_{1,k}(l_{2,k})$ is the number of possible configurations of the QMBS state in the subsystem $B$ with $m_{M/2}=\downarrow(\uparrow)$. The maximum rank of the matrix (79) is two, and the matrix (79) is real symmetric. Therefore, the number of nonzero eigenvalues is at most two. We can show that an eigenvector of the matrix (79) is given by $\bm{x}=[c_{1},c_{1},\ldots,c_{1},c_{2},c_{2},\ldots,c_{2}]^{\rm T}$, where the number of $c_{1}$ and $c_{2}$ are given by $D_{1,k}$ and $D_{2,k}$, respectively. The nonzero eigenvalue can be obtained by calculating the eigenvalues of the following $2\times 2$ matrix:
$\frac{1}{N(M,n)}\begin{bmatrix}l_{1,k}D_{1,k}&l_{2,k}D_{2,k}\\ l_{2,k}D_{1,k}&l_{2,k}D_{2,k}\end{bmatrix}.$ (84)
The eigenvalues are given by
$\lambda_{k,\pm}$ $=\frac{1}{2N(M,n)}\left[l_{1,k}D_{1,k}+l_{2,k}D_{2,k}\pm\sqrt{(l_{1,k}D_{1,k}-l_{2,k}D_{2,k})^{2}+4D_{1,k}D_{2,k}l_{2,k}^{2}}\right].$ (85)
The von Neumann EE can be written as
$S_{\rm vN}$ $=-\sum_{k=0}^{K}\sum_{\eta=\pm}\lambda_{k,\eta}\ln\lambda_{k,\eta}.$ (86)
We can easily evaluate Eq. (86) numerically, even for large system sizes.
Appendix F Edge state in the Krylov space
In this appendix, we discuss the origin of the slow relaxation dynamics starting from the xNéel state. To investigate this, we consider the basis in the Krylov subspace, which we define as $\mathcal{K}(\hat{H},\left|\psi_{0}\right\rangle)\equiv{\rm span}(\left|\psi_{0}\right\rangle,\hat{H}\left|\psi_{0}\right\rangle,\hat{H}^{2}\left|\psi_{0}\right\rangle,\ldots)$, where $\hat{H}$ is a hermitian operator and $\left|\psi_{0}\right\rangle$ is a root state. We assume that the root state is normalized to unity: $\left\langle\psi_{0}|\psi_{0}\right\rangle=1$. Using the standard Lanczos algorithm, we can construct the orthonormal basis $\left|u_{i}\right\rangle$ in the Krylov subspace:
$\hat{H}\left|u_{j}\right\rangle$ $=\alpha_{j}\left|u_{j}\right\rangle+\beta_{j-1}\left|u_{j-1}\right\rangle+\beta_{j}\left|u_{j+1}\right\rangle,\quad(j=1,2,\cdots,N_{\rm K})$ (87)
$\alpha_{j}$ $=\left\langle u_{j}\right|\hat{H}\left|u_{j}\right\rangle,$ (88)
$\beta_{j}$ $=\sqrt{\left[\left\langle u_{j}\right|(\hat{H}-\alpha_{j})-\beta_{j-1}\left\langle u_{j-1}\right|\right]\left[(\hat{H}-\alpha_{j})\left|u_{j}\right\rangle-\beta_{j-1}\left|u_{j-1}\right\rangle\right]},$ (89)
where $\left|u_{0}\right\rangle=0$, $\left|u_{1}\right\rangle=\left|\psi_{0}\right\rangle$, $\beta_{0}=0$, $\beta_{N}=0$, and $N_{\rm K}$ is the dimension of the Krylov subspace. Here, $\alpha_{j}$ and $\beta_{j}$ are real by definition. The matrix representation of the Hamiltonian in the Krylov basis becomes the tridiagonal form:
$\left\langle u_{i}\right|\hat{H}\left|u_{j}\right\rangle$ $=\begin{bmatrix}\alpha_{1}&\beta_{1}&&&\\ \beta_{1}&\alpha_{2}&\beta_{2}&&\\ &\beta_{2}&\alpha_{3}&\beta_{3}&&\\ &&\ddots&\ddots&\ddots&\\ &&&\beta_{N_{\rm K}-2}&\alpha_{N_{\rm K}-1}&\beta_{N_{\rm K}-1}\\ &&&&\beta_{N_{\rm K}-1}&\alpha_{N_{\rm K}}\end{bmatrix}.$ (90)
[…]
{^$\{\hat{\mathcal{I}}\hat{\mathcal{T}},\hat{H}_{\rm DH}\}=0,$ (91)
[…]
From Eqs. (91) and (92), we obtain
[…]
For $k=1$, this relation is equivalent to $\alpha_{1}=0$. From $\alpha_{1}=0$, it follows that $\alpha_{2}=0$:
$\alpha_{2}=\left\langle u_{2}\right|\hat{H}_{\rm DH}\left|u_{2}\right\rangle$ $=\frac{1}{\beta_{1}^{2}}\left\langle u_{1}\right|\hat{H}^{3}_{\rm DH}\left|u_{1}\right\rangle=0.$ (94)
We can also show that $\left|u_{2}\right\rangle$ is an eigenstate of ^$\hat{\mathcal{I}}\hat{\mathcal{T}}$ and has the opposite parity to $\left|u_{1}\right\rangle$:
^$\hat{\mathcal{I}}\hat{\mathcal{T}}\left|u_{2}\right\rangle$ $=-(-1)^{M/2}\left|u_{2}\right\rangle.$ (95)
Using these relations, we can show $\alpha_{j}=0$ by induction.
Figure 14: (a) Krylov coefficient $\beta_{i}$ for $M=16$ and $h^{x}=0.1D$. (b) Magnified view of (a).
From the above results, we can write down the matrix representation of the DH Hamiltonian as
$H_{\rm DH}$ $=D\begin{bmatrix}0&\beta_{1}&&&\\ \beta_{1}&0&\beta_{2}&&\\ &\beta_{2}&0&\beta_{3}&&\\ &&\ddots&\ddots&\ddots&\\ &&&\beta_{N_{\rm K}-2}&0&\beta_{N_{\rm K}-1}\\ &&&&\beta_{N_{\rm K}-1}&0\end{bmatrix}.$ (96)
We can regard this matrix as a one-dimensional tight-binding Hamiltonian with nonuniform nearest-neighbor hopping amplitude $\beta_{j}$. We can easily find that the unitary matrix $\Gamma={\rm diag}(1,-1,1,-1,\ldots)$ satisfies the relation $\{\Gamma,H_{\rm DH}\}=0$. Because all the matrix elements of the matrix $H_{\rm DH}$ are real, we find that this Hamiltonian belongs to the symmetry class BDI in the Altland-Zirnbauer classification Altland1997 . According to the periodic table of topological insulators and superconductors, edge states are allowed in one-dimensional class BDI systems Schnyder2008 .
Here, we numerically diagonalize the Hamiltonian (96). The $j$th eigenvector of the DH Hamiltonian is denoted by $\left|\Psi^{(j)}\right\rangle=\sum_{i}c_{i}^{(j)}\left|u_{i}\right\rangle$. We plot the overlap between each eigenstate and the xNéel state (or $\left|u_{1}\right\rangle$) in Fig. 15 (a). We can find an almost zero-energy state with a large overlap with the xNéel state. We plot the site dependence of the state in Fig. 15 (b). This result clearly shows that the zero-energy eigenstate is localized around $j=1$. We can regard this localized state as a topological edge state. Therefore, we conclude that the slow relaxation dynamics from the xNéel state can be regarded as a consequence of the topological edge states in the Krylov space. We find similar discussions in the context of the Krylov complexity Yates2020 ; Yates2020_2 ; Rabinovici2021 ; Rabinovici2022 ; Trigueros2022 ; Bhattacharjee2022 .
Figure 15: (a) Overlap between the xNéel state and eigenstate of the Hamiltonian (96) for $M=16$ and $h^{x}=0.1D$. The blue square represents the state with the largest overlap with the xNéel state. (b) The site dependence of the wave function of the zero-energy eigenstate is marked by the blue square in (a).